# Optimal Average Disk-Inspection via Fermat’s Principle

### v2

@@ 6 Discussion
Bellman introduced the famous lost-in-a-forest problem and proposed several variants almost seventy years ago [9]. In this work we resolve one of these variants, the Average-Case Disk-Inspection problem. The line of inquiry began with the heuristics of Gluss [28] in the 1960s and continued through the discretization framework developed recently in [15, 16]. Our analysis not only establishes the exact optimum with certified numerical accuracy, but also reveals the structural nature of optimal trajectories. We show that they arise from a reformulation of the problem as an optics model based on Fermat’s Principle of Least Time, which leads to a single-parameter ODE system. Crucially, the resulting optimal trajectories avoid the unit disk, contrary to the conjecture of Gluss.
@@ Acknowledgements
Special thanks to Derek Muller and his team, whose YouTube channel Veritasium and in particular the episode “The Closest We’ve Come to a Theory of Everything” (https://www.youtube.com/watch?v=Q10_srZ-pbs) inspired the main ideas behind this work. Special thanks also to Caleb Jones for early discussions of the project.
@@
This integral form admits a unique continuous solution near $x=0$ by a contraction mapping argument. Expanding once gives $\varepsilon(x)=\pi x+O(x^{3})$, hence
@@
Set $h_{i}=i/n$ and $\Delta h=1/n$, with $\alpha=2\pi/n$. Define the piecewise linear interpolant $\tau_{n}:[0,1]\to\mathbb{R}$ by $\tau_{n}(h_{i})=t_{i}$ for $i=0,\ldots,n$. By Definition 2.5 and Lemma 5.1, we have $\tau_{n}\to\tau$ uniformly on compact subsets of $(0,1]$ and $\tau(0)=\tau_{0}$.
@@
$A_{i}=L_{i}(t_{i})=L(h_{i},t_{i})=\left(\begin{array}[]{c}\cos(2\pi h_{i})-t_{i}\sin(2\pi h_{i})\\ -\sin(2\pi h_{i})-t_{i}\cos(2\pi h_{i})\end{array}\right).$
By Definition 2.5, the curve $\mathcal{T}$ satisfies $\mathcal{T}(x)=L(x,\tau(x))$ for all $x\in[0,1]$. Hence
@@
Since $\tau_{n}(h_{i})=t_{i}$ and $\tau_{n}\to\tau$ uniformly on compact subsets of $(0,1]$, it follows that $\max_{i:h_{i}\in[\delta,1]}\lVert A_{i}-\mathcal{T}(h_{i})\rVert\to 0$ for every $\delta>0$.
Let $\widetilde{\mathcal{T}}_{n}$ be the polygonal path obtained by linear interpolation along the segments $A_{i-1}A_{i}$ on each interval $h_{i-1},h_{i}]$. The map $(x,t)\mapsto L(x,t)$ is continuous on $0,1]\times K$ for any compact $K\subset\mathbb{R}$ containing the ranges of $\{t_{i}\}$ and $\tau$. Using uniform continuity of $L$ on $\delta,1]\times K$ together with the convergence at the grid points gives
$\sup_{x\in[\delta,1]}\lVert\widetilde{\mathcal{T}}_{n}(x)-\mathcal{T}(x)\rVert\to 0\quad\text{as }n\to\infty,$
that is, the polygonal trajectory converges uniformly on compact subsets of $(0,1]$ to $\mathcal{T}$. Finally,

### v3

@@ 6 Discussion
Bellman introduced the famous lost-in-a-forest problem and proposed several variants almost seventy years ago [bellman1956minimization]. In this work we resolve one of these variants, the Average-Case Disk-Inspection problem. The line of inquiry began with the heuristics of Gluss [gluss1961alternative] in the 1960s and continued through the discretization framework developed recently in [conley2025multiagentdiskinspectionArxiv, conley2025multi]. Our analysis not only establishes the exact optimum with certified numerical accuracy, but also reveals the structural nature of optimal trajectories. We show that they arise from a reformulation of the problem as an optics model based on Fermat’s Principle of Least Time, which leads to a single-parameter ODE system. Crucially, the resulting optimal trajectories avoid the unit disk, contrary to the conjecture of Gluss.
@@ Acknowledgements
We thank Derek Muller and his team at Veritasium for an expository discussion of Fermat’s principle of least time, in particular the episode “The Closest We’ve Come to a Theory of Everything”, without which the perspective adopted in this work would likely not have emerged. Special thanks also to Caleb Jones for early discussions of the project.
@@
We show next that this integral form admits a unique continuous solution near $x=0$ by a contraction mapping argument. Indeed, integrating from 0 to $x$ gives the fixed point form
$\varepsilon(x)=\pi x-\frac{1}{x}\int_{0}^{x}r(\varepsilon(u))\,\operatorname{d}\mkern-2.0mu{u}\qquad(x\in(0,\delta]).$
Fix $a>0$ and consider the complete metric space
$X=\{f\in C([0,\delta]):f(0)=0\text{ and }\|f\|_{\infty}\leq a\}$
with the sup norm. Define $T:X\to C([0,\delta])$ by $(Tf)(0)=0$ and $(Tf)(x)=\pi x-\frac{1}{x}\int_{0}^{x}r(f(u))\,\operatorname{d}\mkern-2.0mu{u}\quad(x\in(0,\delta]).$ Since $r(\varepsilon)=\tan\varepsilon-\varepsilon$ is $C^{1}$ near 0 with $r(0)=r^{\prime}(0)=0$, there exists $C>0$ such that for all $|y|,|z|\leq a$,
$|r(y)|\leq C|y|^{3}\quad\text{and}\quad|r(y)-r(z)|\leq Ca^{2}|y-z|.$
Choosing $a>0$ so that $Ca^{2}\leq\tfrac{1}{2}$ and $\delta\leq\tfrac{a}{2\pi}$ ensures that $T(X)\subseteq X$ and that $T$ is a contraction on $X$. By the Banach fixed point theorem, $T$ has a unique fixed point $\varepsilon\in X$, yielding a unique continuous solution near $x=0$.
Expanding the fixed point equation once gives $\varepsilon(x)=\pi x+O(x^{3})$, hence
@@
Set $h_{i}=i/n$ and $\Delta h=1/n$, with $\alpha=2\pi/n$. Define the piecewise linear interpolant $\tau_{n}:[0,1]\to\mathbb{R}$ by $\tau_{n}(h_{i})=t_{i}$ for $i=0,\ldots,n$.
Fix $0<\delta\leq\xi^{\prime}<\xi$. By Definition 2.5 and Lemma 5.1, we have $\tau_{n}\to\tau$ uniformly on $\delta,\xi^{\prime}]$ and $\tau(0)=\tau_{0}$.
@@
$A_{i}=L_{i}(t_{i})=L(h_{i},t_{i}).$
By Definition 2.6, the curve $\mathcal{T}$ satisfies $\mathcal{T}(x)=L(x,\tau(x))$ for all $x\in[0,1]$. Hence, for every $i$ with $h_{i}\in[\delta,\xi^{\prime}]$,
@@
Since $\tau_{n}(h_{i})=t_{i}$ and $\tau_{n}\to\tau$ uniformly on $\delta,\xi^{\prime}]$, it follows that
$\max_{i:~h_{i}\in[\delta,\xi^{\prime}]}\lVert A_{i}-\mathcal{T}(h_{i})\rVert\to 0.$
Let $\widetilde{\mathcal{T}}_{n}$ be the polygonal path obtained by linear interpolation along the segments $A_{i-1}A_{i}$ on each interval $h_{i-1},h_{i}]$. Let $K\subset\mathbb{R}$ be a compact interval containing $\tau([\delta,\xi^{\prime}])$ and containing $\tau_{n}([\delta,\xi^{\prime}])$ for all sufficiently large $n$ (this exists by uniform convergence). The map $(x,t)\mapsto L(x,t)$ is uniformly continuous on $\delta,\xi^{\prime}]\times K$. Combining uniform continuity of $L$ on $\delta,\xi^{\prime}]\times K$ with the convergence at the grid points yields
$\sup_{x\in[\delta,\xi^{\prime}]}\lVert\widetilde{\mathcal{T}}_{n}(x)-\mathcal{T}(x)\rVert\to 0\quad\text{as }n\to\infty.$
Since $0<\delta\leq\xi^{\prime}<\xi$ was arbitrary, this proves the claimed convergence on $0,\xi)$.
Finally,